alright here goes lol. i tryed to solve it and came up with some strange answers.
1. 4x+2y+z=7
2x+2y-4z=-4
x+3y-2z=-8
My attempts: (sorry for the confusing factor x.x)
(4x+2y+z=7)4
2x+2y-4z=28
_____________
2x+2y=-4
16x+8y+28
___________
18x+10y=24
(4x+2y+z=7)2
x+3y-2z=-8
____________
8x+4x+2z=7
(9x+7y=-1)-2
18x+10y=24
____________
-18x-14y=2
18x+10y=24
___________
-4y = 26
___ ___
-4 -4
y=-6.5
18x+10(-6.5)=24
18x-65=24
__________
18x=89
x=4.94
4(4.94)+2(-6.5)+z=7
19.70-13+z=7
6.76+z=7
z=.24
thats it. im sorry its so confusing. im so very confused about this stuff so please help x.x

Found 3 solutions by drk, richwmiller, stanbon:
Answer by drk(1908)

Lets take this one step at a time. Here is the original problem:
(i) 4x+2y+z=7
(ii) 2x+2y-4z=-4
(iii) x+3y-2z=-8
---
basically, we want to kill a variable. getting rid of X look pretty easy
step 1 -> ROW 1 - 2 x ROW 2 AND ROW 2 - 2 x ROW 3 to get
(iv) -2y + 9z = 15
(v) -4y = 12
Luck y us, we can get y right away as
y+=+-3
step 2 -> put the value of y back into (iv) to get y as
-2(-3) + 9z = 15
9z = 9
z=1
Put the values of z and y into (i) to get x as
4x + 2(-3) + (1) = 7
4x -6 + 1 = 7
4x - 5 = 7
4x = 12
x+=+3
so (x,y,z) = (3, -3, 1)

Answer by richwmiller(17219)

x = 3, y = -3, z = 1
Now for your attempt
(4x+2y+z=7)4
2x+2y-4z=28
_____________
2x+2y=-4
16x+8y+28

(4x+2y+z=7) *4
16x+8y+4z=28
What happened to the 4z and why only three terms?
From here on it is trouble.
Let's try again

4x+2y+z=7 (1)
2x+2y-4z=-4 (2)
x+3y-2z=-8 (3)

Answer by stanbon(75887)

4x+2y+z=7
2x+2y-4z=-4
x+3y-2z=-8
-------------------
Rearrange the equations
x+3y-2z=-8
2x+2y-4z=-4
4x+2y+z=7
-----------------------
Divide thru the 2nd equation by 2:
x+3y-2z=-8
x+y-2z=-2
4x+2y+z=7
--------------------
Subtract the 1st from the 2nd
Subtract 4 times the 2nd from the 3rd
x+3y-2z=-8
0x-2y+0z=6
0x-2y+9z=15
-------------------
Solve the 2nd equation for "y":
y = -3
-------------------
Solve the 3rd 3equation for "z":
-2*-3 + 9z = 15
9z = 9
z=1
------------------------------------
Substitute for y and for z to solve for "x":
x + 3*-3 -2*1 = -8
x - 11 = -8
x = 3
==========================
x = 3 ; y = -3 ; z = 1
==============================
Cheers,
Stan H.
==============================